On April 1, the unpaid balance in an account was $218. A payment of $30 was made on April 11. On April 21, a $40 purchase was made. The finance charge rate was 15% per month of the average daily balance. Find the new balance at the end of April.
Answer by Theo(13342)

on april 1, the unpaid balance was 218.
on april 11, a payment of 30 was made.
on april 21, a 40 purchase was made.

from april 1 to april 10, the unpaid balance was 218.
from april 11 to april 20, the unpaid balance was 218 - 30 = 188 because a 30 dollar deposit to the account was made.
from april april 21 to april 30, the unpaid balance was 188 + 40 = 228 because a 40 dollar purchase was made.

you want to find the total daily balance for the month and then divide by the number of days in the month.

the total balance was:

april 1 to april 10 is 10 days * 218 = 2180.
april 11 to april 20 is 10 days * 188 = 1880.
april 21 to april 30 is 10 days * 228 = 2280.

the total is 6340.

divide that by 30 to get an average daily balance of 211.3333333.....

the finance charge would be .15 * 211.33333..... = 31.7

here's a tutorial that can help you to understand.

http://college-cram.com/study/finance/interest/average-daily-balance/

here's another one.

http://credit.about.com/od/creditcardbasics/qt/avgdailybalance.htm

keep in mind that the method is probably the same, but different banks might calculate it differently.

some will use the actual days in the month.
some will use a standard 30 day month.
the bottom line is that they don't all make the same assumptions, so check the bank you're working with to see what their policy is.